OpenStudy (anonymous):

Applications of Solving Systems of Equations Find the two numbers whose sum is 51 and whose difference is 13. show work and check step please

OpenStudy (anonymous):

x+y=51 x-y=13 Add both equations: x+x+y+(-y)=51+13, thus:2x=64, then x=32 and y=32-13=19 Check: 32+19=51 and 32-19=13

OpenStudy (campbell_st):

set up the equations 1st number x 2nd number y x + y = 51 x - y = 13 add the 2 equations to eliminate y 2x = 64 I'll let you finish it... when you have x, you should be able to find y by substituting into either eqauation

OpenStudy (anonymous):

thank you all